Electric field at point p(2m,1m,2m) due to a point charge 5.0 muC placed at origin is


1
Expert's answer
2021-05-04T12:06:43-0400

Find the distance squared from the origin to the point:


"D^2={2^2+1^2+2^2}=9\text{ m}^2."

The field at this point:


"E=\frac{kQ}{D^2}=4993\text{ V/m}."